# Test time training enhances in-context learning of nonlinear functions

### v1

@@ Appendix D One step gradient descent for weak recovery
From Lemma 8, we have that
@@
where $P_{1}^{\prime}=o_{d}(1)$, $P_{2}^{\prime}=\Theta_{d}((\log{d})^{-C_{P_{2}}})$ and $n_{3}=o_{d}(P_{2}^{\prime}r^{-\mathrm{ge}(\sigma_{*})/2-1/2}\log^{-2\deg(\sigma_{*})+2}d)$. Noticing that $g(\Gamma^{*},{\bm{x}})=O(1)$ with high probability, then, letting $b=\frac{1}{N_{new}}\sum_{i=1}^{N_{new}}g(\Gamma^{*},{\bm{w}}_{i})$ and applying Lemma 26, it holds that
@@
Following the same argument as Lemma 6 yields
@@
with high probability. Next we investigate the effect of $n_{3}$. From Lemma 9, $\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}mC_{u})=\tilde{O}(\alpha_{1}mr^{-1/2})$ holds. This leads to $n_{3}\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}mr^{-\mathrm{ge}(\sigma_{*})/2-1})$. Moreover, from Lemma 11, when $\mathrm{ge}(\sigma_{*})=2$ we have $\gamma({\bm{x}},y)=O(1/\sqrt{r})$. Combining this fact with $N_{1}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*})+1})$ yields $\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}m/r)$. Therefore, whether $\mathrm{ge}(\sigma_{*})=1$ or $\mathrm{ge}(\sigma_{*})=2$, we obtain
@@
Lemma 13.
Under the condition of Lemma 12, it holds that
@@
As $\|\beta\|=1$, using the result of Lemma 12 and taking the inner product of $\bm{h}$} and $\beta$ yields
@@
We notice that the leading term of the right hand side is $n_{3}\alpha_{1}L_{m}P_{0}\left\langle{\bm{u}},{\bm{x}}\right\rangle$ when $\mathrm{ge}(\sigma_{*})=1$, and $n_{3}\alpha_{1}mP_{1}\left\langle\beta,\sqrt{r}\Gamma_{*}^{\top}\Gamma_{u}{\bm{x}}\right\rangle/\rho\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle$ when $ge_{(}\sigma_{*})=2$. Note that $\mathbb{E}_{{\bm{x}}}[P_{0}\left\langle{\bm{u}},{\bm{x}}\right\rangle]=0$ and $\mathbb{E}_{{\bm{x}}}[P_{1}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle/\rho\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle]=P_{1}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{u}}\right\rangle/\rho=\tilde{O}(r^{-3/2})$ because of $\sqrt{r}\Gamma_{u}^{\top}\Gamma^{*}\beta=\tilde{O}(1/\sqrt{r})\beta+\tilde{O}(1/r)+\left\langle\beta,{\bm{u}}\right\rangle{\bm{u}}$ (see Lemma 32). Moreover, we have $\mathbb{E}_{{\bm{x}}}[P_{0}^{2}\left\langle{\bm{u}},{\bm{x}}\right\rangle^{2}]=\tilde{O}(1)$ and $\mathbb{E}_{{\bm{x}}}[P_{1}^{2}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle^{2}/\rho^{2}\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle^{2}]\leq P_{1}^{2}/\rho^{2}\cdot\mathbb{E}[\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle^{4}]^{1/2}\mathbb{E}[\left\langle{\bm{x}},{\bm{u}}\right\rangle^{4}]^{1/2}=\tilde{O}(1)$. Therefore, Lemma 26 yields that
@@ Appendix E Strong recovery
Remember that, in the stage of strong recovery, the initialization scale is $\alpha_{2},$ the context length is $N_{2}$ and $C_{u}=1$. By using the same argument to derive Lemma 6, we have
@@
Now, from Lemma 15, we also have
@@
for any $1\leq j,l\leq N_{3}$ with high probability, where $C=\tilde{O}(1)$. Let $c_{k}=1-\sqrt{\frac{k+1}{k+2}}$. Note that $c_{k}$ is a constant that only depends on $k$, that is $c_{k}=O(1)$. Suppose that $c_{1}\leq a^{(i)}\leq 1-\frac{1}{3}c_{k}$ for all $i=1,2,\dots,N_{3}$. Then, from Lemma 16, we have
@@
In Lemma 9, we assumed that $\left\langle{\bm{u}},{\bm{x}}_{i}\right\rangle=\tilde{O}_{d}(1)$. Now ${\bm{u}}^{(n)}$ and ${\bm{x}}_{i}$ are not independent of each other, so we need to ensure that $\left\langle{\bm{u}}^{(j)},{\bm{x}}_{i}\right\rangle=\tilde{O}_{d}(1)$ for all $j=1,2,\dots,N_{3}+1$.
Lemma 18.
@@
From Lemma 15, we have
@@
Following the same argument as Lemma 15 yields $\left\langle{\bm{x}}_{i}^{*},\nu\right\rangle=\tilde{O}(\alpha_{2}m\sqrt{r})$. Thus, by letting $\nu_{*}=\left\langle{\bm{x}}_{i}^{*},\nu\right\rangle$, we obtain the assertion. ∎
Lemma 19.
Set $\alpha_{2}m=\tilde{\Theta}(\varepsilon/r)$, $N_{2}=\tilde{\Theta}(r^{2})$, $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$ and $\eta_{2}=\tilde{\Theta}(1/\sqrt{r})$(This is exactly the same situation as Lemma 16 and Lemma 17). Then we have
@@
When $j=1$, ${\bm{x}}_{i}$ and ${\bm{u}}^{(1)}$ are independent of each other, thus Lemma 29 yields the desired result. Substituting the parameters into the result of Lemma 18 yields
@@
This lemma allows us to use Lemma 9. Note that $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$ means, by ignoring the term $\log{\frac{1}{\varepsilon}}$, $\varepsilon=\tilde{\Theta}({\frac{r\sqrt{r}}{N_{3}}})$. Now we can estimate $\left\langle\beta,{\bm{x}}\right\rangle$ using $\left\langle{\bm{u}}^{(N_{3}+1)},{\bm{x}}\right\rangle$. For notational simplicity, we write ${\bm{u}}^{(N_{3}+1)}$ as $\bm{u}$} from now on.
Lemma 20.
@@
Then, combining with Lemma 25 yields $\left\langle\beta-{\bm{u}},{\bm{x}}\right\rangle=O_{p}(\sqrt{2\varepsilon\log{d}})$. ∎
@@ Appendix F Training MLP layer
Lemma 22.
Under the condition of Lemma 21, there exists ${\bm{a}}^{\prime}\in\mathbb{R}^{m}$ such that
@@
Using $\pi(v,b)$ in Lemma 21, define $a_{j}^{\prime}=m^{-1}\pi(v_{j},b_{j})$. Since $\sup_{v,b}|\pi(v,b)|=\tilde{O}(1)$, from Hoeffdng’s inequality, it holds that
@@
Let $M=N_{1}+N_{2}+N_{3}$ and $\bm{a}$$^{\prime}$ be the output parameter constructed in Lemma 22: from the equivalence between $\ell_{2}$-regularized and norm-constrained optimization algorithms, if we carefully choose $\lambda_{3}$, then we have
@@
(recall that we assumed $\tau=\Theta(1)$) with high probability, which yields the first assertion. Moreover, from Lemma 22, we can see that
@@
Appendix G Proof of the Theorem 1
Now, we are ready to give the proof of Theorem 1.
@@ Appendix G Proof of the Theorem 1
From Lemma 23,the first two terms $\frac{1}{N_{4}}\sum_{i=N_{1}+N_{2}+N_{3}+1}^{N_{!}+N_{2}+N_{3}+N_{4}}\left|y_{i}-f_{\mathrm{TF}}({\bm{x}}_{i},{\bm{u}},{\bm{v}}^{*},{\bm{a}}^{*},{\bm{b}}^{*})\right|-\tau$ are bounded by $\tilde{O}(m^{-1/2})+O(\delta\log^{2\deg(\sigma_{*})-2}d)$. Moreover, from Lemma 22 and the definition of $\bm{b}$$^{*}$, we have $\|{\bm{a}}^{*}\|=\tilde{O}(m^{-1/2})$ and $\|{\bm{b}}^{*}\|=\tilde{O}(\sqrt{m})$. Therefore, from Lemma 24, we have
@@
Finally, we rewrite the term $O(\delta\log^{2\deg(\sigma_{*})-2}d)$. From Lemma 20, $\delta=O(\sqrt{2\varepsilon\log{d}})$ holds. Also, $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$, which means, when we ignore the term $\log{\frac{1}{\varepsilon}}$, $\varepsilon=\tilde{\Theta}({\frac{r\sqrt{r}}{N_{3}}})$. Therefore
@@ Appendix H Additional Lemmas
Lemma 27 (Damian et al. (2023), Property 1).
@@ H.1 Pretrained matrix
From lemma 31, we have $\sqrt{r}\kappa\Gamma^{*}\beta=\beta+{\bm{d}}$ where ${\bm{d}}=O(1/\sqrt{d})$. Therefore, it holds that
@@
Again from Lemma 31, we see that $\sqrt{r}\kappa\Gamma^{*}\beta=\beta+O(1/\sqrt{d})$. Also we have $\sqrt{r}\kappa\Gamma^{*}{\bm{d}}={\bm{U}}^{\top}{\bm{D}}{\bm{U}}{\bm{d}}+{\bm{N}}{\bm{d}}$, and since $\|{\bm{U}}^{\top}{\bm{D}}{\bm{U}}\|_{2}=1$ from the definition of $\bm{U}$} and $\bm{D}$}, we have ${\bm{U}}^{\top}{\bm{D}}{\bm{U}}{\bm{d}}=O(1/\sqrt{d})$. Finally, since $\|{\bm{N}}\|_{F}=O(1/\sqrt{d})$, we have ${\bm{N}}{\bm{d}}=O(1/d)$. This indicates that $\sqrt{r}\kappa\Gamma^{*}{\bm{d}}=O(1/\sqrt{d})$. ∎
@@ Appendix I Experimental details
where $\tilde{{\bm{y}}_{i}}=[y_{i},0,\dots,0]^{\top}$. Next, the read-in layer transforms this embedding into $\tilde{{\bm{E}}}\in\mathbb{R}^{D\times(2N+2)}$, where $D=128$. This mapped embedding $\tilde{{\bm{E}}}$ goes through a 6-layer GPT-2 backbone with 4 attention heads, following the configuration by Garg et al. (2023). Finally, the output of GPT-2 backbone is transformed by the read-out layer into the vector $z_{1},z_{2},\dots,z_{2N+1},z_{2N+2}]$. Here, $z_{2i-1}$ is the prediction of $y_{i}$ given the context $({\bm{x}}_{1},y_{1},\dots,{\bm{x}}_{i-1},y_{i-1},{\bm{x}}_{i})$. We used Adam optimizer (Kingma & Ba, 2017) with a learning rate of 0.0001. To reduce the pretraining cost, we adopted the curriculum learning strategy, which is also used in Garg et al. (2023). The training started with the dimension $d=4$, and the dimension was increased by two until it reached the target dimension.
For the model with TTT, we used the same base model as ICL evaluation. We introduced low-rank adaptation (LoRA) to the attention projection layers (c_attn and c_proj) and the feedforward layer (c_fc). The rank of LoRA was set to $4$, and the parameters LoRA_alpha and LoRA_dropout were set to $8$ and $0.1$, respectively. The LoRA parameters were updated $10$ times for each query. The inference-time learning rate was $0.01$ and $0.05$ in the experiment in Figure 1(a) and Figure 1(b), respectively. To prevent information leakage when evaluating the prediction of $y_{i}$, the model used only the preceding data $({\bm{x}}_{1},y_{1},\dots,{\bm{x}}_{k-1},y_{k-1},{\bm{x}}_{k})$ to update the weight. The LoRA weights were reset before proceeding to the prediction of the next target, $y_{i+1}$.

### v2

@@ Appendix D One step gradient descent for weak recovery
From Lemma B.1, we have that
@@
where $P_{1}^{\prime}=o_{d}(1)$, $P_{2}^{\prime}=\Theta_{d}((\log{d})^{-C_{P_{2}}})$ and $n_{3}=o_{d}(P_{2}^{\prime}r^{-\mathrm{ge}(\sigma_{*})/2-1/2}\log^{-2\deg(\sigma_{*})+2}d)$. Noticing that $g(\Gamma^{*},{\bm{x}})=O(1)$ with high probability, then, letting $b=\frac{1}{N_{new}}\sum_{i=1}^{N_{new}}g(\Gamma^{*},{\bm{w}}_{i})$ and applying Lemma H.2, it holds that
@@
Following the same argument as Lemma 3.2 yields
@@
with high probability. Next we investigate the effect of $n_{3}$. From Lemma C.1, $\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}mC_{u})=\tilde{O}(\alpha_{1}mr^{-1/2})$ holds. This leads to $n_{3}\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}mr^{-\mathrm{ge}(\sigma_{*})/2-1})$. Moreover, from Lemma C.3, when $\mathrm{ge}(\sigma_{*})=2$ we have $\gamma({\bm{x}},y)=O(1/\sqrt{r})$. Combining this fact with $N_{1}=\tilde{\Omega}(r^{\mathrm{ge}(\sigma_{*})+1})$ yields $\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{w}}_{i})=\tilde{O}(\alpha_{1}m/r)$. Therefore, whether $\mathrm{ge}(\sigma_{*})=1$ or $\mathrm{ge}(\sigma_{*})=2$, we obtain
@@
Lemma D.2.
Under the condition of Lemma D.1, it holds that
@@
As $\|\beta\|=1$, using the result of Lemma D.1 and taking the inner product of $\bm{h}$} and $\beta$ yields
@@
We notice that the leading term of the right hand side is $n_{3}\alpha_{1}L_{m}P_{0}\left\langle{\bm{u}},{\bm{x}}\right\rangle$ when $\mathrm{ge}(\sigma_{*})=1$, and $n_{3}\alpha_{1}mP_{1}\left\langle\beta,\sqrt{r}\Gamma_{*}^{\top}\Gamma_{u}{\bm{x}}\right\rangle/\rho\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle$ when $ge_{(}\sigma_{*})=2$. Note that $\mathbb{E}_{{\bm{x}}}[P_{0}\left\langle{\bm{u}},{\bm{x}}\right\rangle]=0$ and $\mathbb{E}_{{\bm{x}}}[P_{1}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle/\rho\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle]=P_{1}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{u}}\right\rangle/\rho=\tilde{O}(r^{-3/2})$ because of $\sqrt{r}\Gamma_{u}^{\top}\Gamma^{*}\beta=\tilde{O}(1/\sqrt{r})\beta+\tilde{O}(1/r)+\left\langle\beta,{\bm{u}}\right\rangle{\bm{u}}$ (see Lemma H.8). Moreover, we have $\mathbb{E}_{{\bm{x}}}[P_{0}^{2}\left\langle{\bm{u}},{\bm{x}}\right\rangle^{2}]=\tilde{O}(1)$ and $\mathbb{E}_{{\bm{x}}}[P_{1}^{2}\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle^{2}/\rho^{2}\cdot\left\langle{\bm{x}},{\bm{u}}\right\rangle^{2}]\leq P_{1}^{2}/\rho^{2}\cdot\mathbb{E}[\left\langle\sqrt{r}\Gamma_{u}^{\top}\Gamma_{*}\beta,{\bm{x}}\right\rangle^{4}]^{1/2}\mathbb{E}[\left\langle{\bm{x}},{\bm{u}}\right\rangle^{4}]^{1/2}=\tilde{O}(1)$. Therefore, Lemma H.2 yields that
@@ Appendix E Strong recovery
Remember that, in the stage of strong recovery, the initialization scale is $\alpha_{2},$ the context length is $N_{2}$ and $C_{u}=1$. By using the same argument to derive Lemma 3.2, we have
@@
Now, from Lemma E.1, we also have
@@
for any $1\leq j,l\leq N_{3}$ with high probability, where $C=\tilde{O}(1)$. Let $c_{k}=1-\sqrt{\frac{k+1}{k+2}}$. Note that $c_{k}$ is a constant that only depends on $k$, that is $c_{k}=O(1)$. Suppose that $c_{1}\leq a^{(i)}\leq 1-\frac{1}{3}c_{k}$ for all $i=1,2,\dots,N_{3}$. Then, from Lemma E.2, we have
@@
In Lemma C.1, we assumed that $\left\langle{\bm{u}},{\bm{x}}_{i}\right\rangle=\tilde{O}_{d}(1)$. Now ${\bm{u}}^{(n)}$ and ${\bm{x}}_{i}$ are not independent of each other, so we need to ensure that $\left\langle{\bm{u}}^{(j)},{\bm{x}}_{i}\right\rangle=\tilde{O}_{d}(1)$ for all $j=1,2,\dots,N_{3}+1$.
Lemma E.4.
@@
From Lemma E.1, we have
@@
Following the same argument as Lemma E.1 yields $\left\langle{\bm{x}}_{i}^{*},\nu\right\rangle=\tilde{O}(\alpha_{2}m\sqrt{r})$. Thus, by letting $\nu_{*}=\left\langle{\bm{x}}_{i}^{*},\nu\right\rangle$, we obtain the assertion. ∎
Lemma E.5.
Set $\alpha_{2}m=\tilde{\Theta}(\varepsilon/r)$, $N_{2}=\tilde{\Theta}(r^{2})$, $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$ and $\eta_{2}=\tilde{\Theta}(1/\sqrt{r})$(This is exactly the same situation as Lemma E.2 and Lemma E.3). Then we have
@@
When $j=1$, ${\bm{x}}_{i}$ and ${\bm{u}}^{(1)}$ are independent of each other, thus Lemma H.5 yields the desired result. Substituting the parameters into the result of Lemma E.4 yields
@@
This lemma allows us to use Lemma C.1. Note that $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$ means, by ignoring the term $\log{\frac{1}{\varepsilon}}$, $\varepsilon=\tilde{\Theta}({\frac{r\sqrt{r}}{N_{3}}})$. Now we can estimate $\left\langle\beta,{\bm{x}}\right\rangle$ using $\left\langle{\bm{u}}^{(N_{3}+1)},{\bm{x}}\right\rangle$. For notational simplicity, we write ${\bm{u}}^{(N_{3}+1)}$ as $\bm{u}$} from now on.
Lemma E.6.
@@
Then, combining with Lemma H.1 yields $\left\langle\beta-{\bm{u}},{\bm{x}}\right\rangle=O_{p}(\sqrt{2\varepsilon\log{d}})$. ∎
@@ Appendix F Training MLP layer
Lemma F.2.
Under the condition of Lemma F.1, there exists ${\bm{a}}^{\prime}\in\mathbb{R}^{m}$ such that
@@
Using $\pi(v,b)$ in Lemma F.1, define $a_{j}^{\prime}=m^{-1}\pi(v_{j},b_{j})$. Since $\sup_{v,b}|\pi(v,b)|=\tilde{O}(1)$, from Hoeffdng’s inequality, it holds that
@@
Let $M=N_{1}+N_{2}+N_{3}$ and $\bm{a}$$^{\prime}$ be the output parameter constructed in Lemma F.2: from the equivalence between $\ell_{2}$-regularized and norm-constrained optimization algorithms, if we carefully choose $\lambda_{3}$, then we have
@@
(recall that we assumed $\tau=\Theta(1)$) with high probability, which yields the first assertion. Moreover, from Lemma F.2, we can see that
@@
Appendix G Proof of the Theorem 1.1
Now, we are ready to give the proof of Theorem 1.1.
@@
From Lemma F.3,the first two terms $\frac{1}{N_{4}}\sum_{i=N_{1}+N_{2}+N_{3}+1}^{N_{!}+N_{2}+N_{3}+N_{4}}\left|y_{i}-f_{\mathrm{TF}}({\bm{x}}_{i},{\bm{u}},{\bm{v}}^{*},{\bm{a}}^{*},{\bm{b}}^{*})\right|-\tau$ are bounded by $\tilde{O}(m^{-1/2})+O(\delta\log^{2\deg(\sigma_{*})-2}d)$. Moreover, from Lemma F.2 and the definition of $\bm{b}$$^{*}$, we have $\|{\bm{a}}^{*}\|=\tilde{O}(m^{-1/2})$ and $\|{\bm{b}}^{*}\|=\tilde{O}(\sqrt{m})$. Therefore, from Lemma F.4, we have
@@
Finally, we rewrite the term $O(\delta\log^{2\deg(\sigma_{*})-2}d)$. From Lemma E.6, $\delta=O(\sqrt{2\varepsilon\log{d}})$ holds. Also, $N_{3}=\tilde{\Theta}(\frac{r\sqrt{r}}{\varepsilon}\log{\frac{1}{\varepsilon}})$, which means, when we ignore the term $\log{\frac{1}{\varepsilon}}$, $\varepsilon=\tilde{\Theta}({\frac{r\sqrt{r}}{N_{3}}})$. Therefore
@@ Appendix H Additional Lemmas
Lemma H.3 ((Damian et al., 2023), Property 1).
@@ H.1 Pretrained matrix
From lemma H.7, we have $\sqrt{r}\kappa\Gamma^{*}\beta=\beta+{\bm{d}}$ where ${\bm{d}}=O(1/\sqrt{d})$. Therefore, it holds that
@@
Again from Lemma H.7, we see that $\sqrt{r}\kappa\Gamma^{*}\beta=\beta+O(1/\sqrt{d})$. Also we have $\sqrt{r}\kappa\Gamma^{*}{\bm{d}}={\bm{U}}^{\top}{\bm{D}}{\bm{U}}{\bm{d}}+{\bm{N}}{\bm{d}}$, and since $\|{\bm{U}}^{\top}{\bm{D}}{\bm{U}}\|_{2}=1$ from the definition of $\bm{U}$} and $\bm{D}$}, we have ${\bm{U}}^{\top}{\bm{D}}{\bm{U}}{\bm{d}}=O(1/\sqrt{d})$. Finally, since $\|{\bm{N}}\|_{F}=O(1/\sqrt{d})$, we have ${\bm{N}}{\bm{d}}=O(1/d)$. This indicates that $\sqrt{r}\kappa\Gamma^{*}{\bm{d}}=O(1/\sqrt{d})$. ∎
@@ Appendix I Experimental details
where $\tilde{{\bm{y}}_{i}}=[y_{i},0,\dots,0]^{\top}$. Next, the read-in layer transforms this embedding into $\tilde{{\bm{E}}}\in\mathbb{R}^{D\times(2N+2)}$, where $D=128$. This mapped embedding $\tilde{{\bm{E}}}$ goes through a 2-layer GPT-2 backbone with 4 attention heads, following the configuration by Garg et al. (2023). Finally, the output of GPT-2 backbone is transformed by the read-out layer into the vector $z_{1},z_{2},\dots,z_{2N+1},z_{2N+2}]$. Here, $z_{2i-1}$ is the prediction of $y_{i}$ given the context $({\bm{x}}_{1},y_{1},\dots,{\bm{x}}_{i-1},y_{i-1},{\bm{x}}_{i})$. We used Adam optimizer (Kingma and Ba, 2017) with a learning rate of 0.0001. To reduce the pretraining cost, we adopted the curriculum learning strategy, which is also used in Garg et al. (2023). The training started with the dimension $d=4$, and the dimension was increased by two until it reached the target dimension.
For the model with TTT, we introduced low-rank adaptation (LoRA) to the attention projection layers (c_attn and c_proj) and the feedforward layer (c_fc) of the base model. The rank of LoRA was set to $4$, and the parameters LoRA_alpha and LoRA_dropout were set to $8$ and $0.1$, respectively. The LoRA parameters were updated $300$ times via SGD for each query. The inference-time learning rate was $0.1$ in the experiment in Figure 1 and Figure 2. To prevent information leakage when evaluating the prediction of $y_{i}$, the model used only the preceding data $({\bm{x}}_{1},y_{1},\dots,{\bm{x}}_{i-1},y_{i-1},{\bm{x}}_{i})$ to update the weight. The LoRA weights were reset before proceeding to the prediction of the next target, $y_{i+1}$. We observed numerical instability (gradient explosion) in a negligible fraction of trials; these instances were excluded from the final results to ensure metric stability.